the sides of a triangle are 8, 15 and 17 units. if each side is doubled, by how many square units will the area of the triangle increase?
Answer by checkley77(12844)

If this is a right triangle then:
area=bh/2
area=15*8/2
area=60 square units.
area=30*16/2
area=240 square units.
The area has increased by 240-60=180 square units or a 4 fold increase.